OpenStudy (anonymous):

A circle with the longest chord of 8 has the same areea as an equilateral triangle with side x. What is the height of that triangle. Explanations please - thanks :)

OpenStudy (anonymous):

i know the diam is 8 a = 16*3.14 = 50.24

OpenStudy (anonymous):

so area of triangle is equal ot 50.24, right?

sam (.sam.):

Longest chord is the diameter, so r=4 Area of circle = area of triangle \[\pi r ^{2}=\frac{\sqrt{3}}{4}(abc)^{\frac{2}{3}}\] \[\pi r ^{2}=\frac{\sqrt{3}}{4}(x ^{3})^{\frac{2}{3}}\] \[\pi r ^{2}=\frac{\sqrt{3}}{4}(x ^{2})\] x=10.77 h could be 9.31